OpenStudy (anonymous):

What is the y- intercept of a line with equation y=2/3x+4.

Parth (parthkohli):

\[\rm y = mx + b\]b is the y-intercept.

hartnn (hartnn):

to get y-intercept, just put x=0 in your equation

Parth (parthkohli):

There are two different ways to do this. Yup @hartnn

OpenStudy (baldymcgee6):

I like @ParthKohli 's way.. you can just look at it :D

Parth (parthkohli):

@DustinMHayes Hello, did you figure it out?

OpenStudy (anonymous):

y=(2/3)x+4

Parth (parthkohli):

Yeah, so what is 'b' here?

OpenStudy (anonymous):

4

hartnn (hartnn):

that is correct y-intercept.

OpenStudy (anonymous):

4 or y=(2/3)x+4?

hartnn (hartnn):

y-intercept is a co-ordinate (0,b) where b here is 4